Linda invested her savings in two investment funds. The amount she invested in Fund A was twice as much as the amount she invested in Fund B. Fund A returned a 7% profit and Fund B returned a 3% profit. How much did she invest in Fund A, if the total profit from the two funds together was $1700?
Answer by mananth(16946)

Fund A 3.00% per annum ------------- x
Fund B 7.00% per annum ------------ y= 2x
Interest----- $1,700.00

3.00% x + 7.00% * 2 x = $1,700.00
0.03 x + 0.07 * 2 x = $1,700.00
0.03 x + 0.14 x = $1,700.00
0.17 x = $1,700.00
/ 0.17
x= $10,000.00 --- Fund A
2 x = $20,000.00 Fund B

CHECK
$300.00 + $1,400.00 = $1,700.00
m.ananth@hotmail.ca